Akosua works as the customer service officer at the Agarwal Eye Clinic located in Tesano, a suburb of




the Greater Accra region. The number of requests from male patients for emergency eye treatment that




she receives during any day may be modeled by a poisson distribution with mean of 2.6.Find the




probability that the number of requests from male patients for emergency eye treatment that Akosua




receives on a particular day is;




(i) Exactly 4




(ii)At least 1




(b) Find the probability that the number of requests from male patients for emergency eye




treatment that Akosua receives during a period of 5 days is 15 or fewer





1
Expert's answer
2022-05-13T13:50:33-0400

Let "X=" the number of of requests from male patients: "X\sim Po(\lambda t)."

(a) "\lambda t=2.6(1)=2.6"

(i)

"P(X=4)=\dfrac{e^{-2.6}(2.6)^4}{4!}=0.141422"

(ii)


"P(X\ge 1)=1-P(X=0)"

"=1-\dfrac{e^{-2.6}(2.6)^0}{0!}=0.925726"

(b)


"\lambda t=2.6(5)=13"

"P(X\le 15)=P(X=0)+P(X=1)+P(X=2)"

"+P(X=3)+P(X=4)+P(X=5)"

"+P(X=6)+P(X=7)+P(X=8)"

"+P(X=9)+P(X=10)+P(X=11)"

"+P(X=12)+P(X=13)+P(X=14)"

"+P(X=15)=\dfrac{e^{-13}(13)^0}{0!}+\dfrac{e^{-13}(13)^1}{1!}"

"+\dfrac{e^{-13}(13)^2}{2!}+\dfrac{e^{-13}(13)^3}{3!}+\dfrac{e^{-13}(13)^4}{4!}"

"+\dfrac{e^{-13}(13)^5}{5!}+\dfrac{e^{-13}(13)^6}{6!}+\dfrac{e^{-13}(13)^7}{7!}"

"+\dfrac{e^{-13}(13)^8}{8!}+\dfrac{e^{-13}(13)^9}{9!}+\dfrac{e^{-13}(13)^{10}}{10!}"

"+\dfrac{e^{-13}(13)^{11}}{11!}+\dfrac{e^{-13}(13)^{12}}{12!}+\dfrac{e^{-13}(13)^{13}}{13!}"

"+\dfrac{e^{-13}(13)^{14}}{14!}+\dfrac{e^{-13}(13)^{15}}{15!}=0.76361"